What volume of water should be added to 150mL of 1.10M H2SO4 to obtain a 0.200M solution?
1
Expert's answer
2020-08-02T15:52:18-0400

using the dilution formula M1V1=M2V2

where M is the molarity of the solution and V is the volume

M1= Initial molarity, M2=final molarity, V1=initial volume and V2=final volume

V1=150ml

M1= 1.1M

M2=0.2M

hence, 150*1.1=0.2V2

V2=(150*1.1)/0.2

V2 = 825mL

added volume = 825-150=675mL